OpenStudy (anonymous):

also please destribe which method you would use for solving this ODE: y"+9y=0; y(pie/3)=1, y'(pie/3)=-2

OpenStudy (amistre64):

does seperation of church and state work on doubles?

myininaya (myininaya):

just solve r^2+9=0 to figure out your solution form

OpenStudy (amistre64):

thats what we call an imaginary solution

OpenStudy (amistre64):

...and by we i mean me and my voices

myininaya (myininaya):

mr.paul has the best site i have ever seen for talking about differential equations: http://tutorial.math.lamar.edu/Classes/DE/HOHomogeneousDE.aspx

OpenStudy (amistre64):

dy/dx + 9y = 0 dy/dx = -9y dy/y = 9dx ln(y) = 9x

myininaya (myininaya):

\[y=c_1\cos(3x)+c_2\sin(3x)\]

OpenStudy (amistre64):

i keep sleeping with my head on an ODE book, but i dont feel none to smarter

myininaya (myininaya):

since \[r=\pm 3i \]

OpenStudy (anonymous):

this isnt an exact equation?

OpenStudy (anonymous):

supose a solution of the form: \[y=e ^{ax}\] replace in the equation: \[a ^{2}e ^{ax}+ 9e ^{ax} =0\] thus: \[a ^{2}=-9\] \[a=\pm3i\] This entails: \[y=Ae ^{3ix}+Be ^{-3ix}\]

myininaya (myininaya):

if you have something in the form \[a_{n}y^{(n)}+a_{n-1}y^{(n-1)}+ \cdot \cdot \cdot +a_3y^{(3)}+a_2y''+a_1y'+a_0y=0\] we can find what our solution should be if we solve the polynomial: (well as along as the a's are constant coefficients) \[a_nr^n+a_{n-1}r^{n-1}+\cdot \cdot \cdot a_3r^3+a_2r^2+a_1r+a_0=0\]

myininaya (myininaya):

for instance we had \[y''+9y=0\] this means all we really have to do is solve \[r^2+9=0=>r=\pm 3i \] => \[y=Acos(3x)+Bsin(3x)\] :) if we have \[y''-9y=0\] then solution is \[y=ae^{3x}+be^{-3x}\] since the solutions would have been r=pm 3

OpenStudy (anonymous):

ooo gotcha

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

Refer to the attachment.